A piggyback of 750 mg / 250mlis infusing at 50 ml/hr. How much of the drig has the patient recieved in 4 hours?
Answer #1

The bag contains 750 mg/ 250 ml ( 750 mg of the drug in 250 ml of fluid)

Therefore 1 ml contains - 750/250 = 3mg

  • so 1 ml contains 3 mg of drug ------- (1)
  • The infusion is being administered at the rate of - 50 ml/hr

Now, coming to the question- how much drug has the patient received in 4 hrs?

  • If the infusion was going at the rate of 50ml/hr- then a total of 50ml + 50ml + 50ml + 50ml = 200 ml has been infused over 4 hours ( 50 ml for each hour).
  • from statement (1) we know that 1 ml of infusate contains 3 mg of the drug
  • So 200 ml of the infusate will contain = 3 * 200 = 600 mg

Therefore, the patient has received 600mg of the drug over 4 hour